Calculate the pH at 25°C of a 0.0019 M solution of a weak base with a Kb of 3.2 × 10−9.
Answer #1

Answer:

Let the weak base be represented by B-OH

Given the concentration = 0.0019 M and the equillibrium constant Kb = 3.2 x 10-9

The dissociation equation is given by

               B-OH    B+    +           OH-

The initial concentration       0.0019                         0                            0

                                                 Let x moles dissociated at equillibrium.

equillibrium concentration      0.0019 -x                      x                            x

                  Kb = [B+] [OH-] / [B-OH]

                        = x * x / [0/0019 -x]

                        = x2 / 0.0019          since x is small compared to 0.0019, it is neglected in the denominator.

but kb =   3.2 x 10-9           (given)

Therefore         3.2 x 10-9 = x2 / 0.0019

                           x2 = 3.2 * 10-9 * 0.0019

                                = 6.08 * 10-12

Therefore x = (6.08 x 10-12 )

                 = 2.466 x 10-6

Concentration of [OH] = 2.466 x 10-6

POH = - log [OH-] = - log [2.466 x 10-6]

                             = 6 - log 2.466

                             = 6 - 0.4 = 5.6

We know that PH + POH = 14

                    PH = 14 - POH

                         = 14 - 5.6

                         = 8.4

Thus

The PH of the given weak base solution is 8.4